OpenStudy (anonymous):

Find S15 for the series -1 + -3 + -5 + -7 +...Type the numeric answer only.

OpenStudy (anonymous):

will give medal!

OpenStudy (austinl):

-1-3-5-7-9-11-13-15-17-19-21-23-25-27-29-31=S15

OpenStudy (anonymous):

thanks! Find the 6th term of the sequence in which t1 = -5 and tn = -5tn-1?

OpenStudy (anonymous):

please help!!!

OpenStudy (austinl):

n=6 \[t_{1}=-5\] \[t_{n}=-5t_{n}-1\] Is this correct?

OpenStudy (anonymous):

im not sure??? i dont know how to do this

OpenStudy (amistre64):

no, they give the recurrsion, not the explicit

OpenStudy (amistre64):

\[t_{n}=-5t_{n-1}\]

OpenStudy (austinl):

Ok, that makes much more sense. I wish that people would use proper format.

OpenStudy (amistre64):

it would help yes :)

OpenStudy (austinl):

Ok, so we are given \[t_{1}=5\] \[t_{2}=-5\times t_{1}\]

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

so -5 times 5?

OpenStudy (austinl):

Right! And that is....?

OpenStudy (anonymous):

-25

OpenStudy (austinl):

Correct, so \[t_{2}=-25\] \[t_{3}=-5\times t_{2}\] \[t_{4}=-5\times t_{3}\] \[t_{5}=-5\times t_{4}\] \[t_{6}=-5\times t_{5}\]

OpenStudy (anonymous):

so the ansswer would be -15,625?

OpenStudy (austinl):

That is the answer that I arrived to!

OpenStudy (anonymous):

okay are you willing to help me get started on one more? (:

OpenStudy (austinl):

Sure, post up a new question so that others will be more likely to see it!

OpenStudy (anonymous):

okay!